OpenStudy (anonymous):

Question One: find [goh](x) and [hog](x) g(x)=6x h(x)=-10x^3+9x^2-7x+6 Question Two: What is the inverse of the point (-3,15)

OpenStudy (anonymous):

Determine whether each pair of functions are inverse functions 1. f(x)=7x+3/2 g(x)=11x-3/7 2. f(x)=x-14 g(x)=x+14

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

I figured out the pair of functions question

OpenStudy (anonymous):

I just need question one and question two